a body of mass 500g moves with a speed of 4m/s is a circular path of radious 5m. 
(1) find the centripetal force on the body.
(2) frequency of the body in circle per second

Grade:6

1 Answers

Vikas TU
7 years ago
First:
Centripital Force = mv^2/r
                        = 0.5*16/5
                        = 8/5 N.
Second:
Time Period = 2pir/v => 2*3.14*5/4 => 7.85
Frequency = 1/T => 0.1273888